40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 194 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  7328   Fri Aug 31 13:22:29 2012 ElliUpdateIOOCorrecting improper termination of the 55MHz input to the EOM

Koji, Riju, Elli

This morning Koji discovered that the 55MHz input into the RF combiner that I disconnected yesterday wasn't terminated properly, so it was reflecting power back into the amplifier in the signal generation unit.    We turned off the signal generation unit and checked that the amplifier was still working properly- it was.  A 50 ohm terminator was attached to the end of the 55MHz cable so that it is now terminated properly.

When we tried to turn the signal generator box back on we discovered the switch is broken (the box will only stay on while you hold down the on switch) and will need to be replaced.  In order to create the 29.5MHz sidebands to lock the mode cleaner, we bypassed the signal generation unit which won't stay on (unplugging '29.5 MHZ out' cable from the frequency generation unit), and instead sent a 0.39V 29.5MHz signal from a function generatior into 'RF input' on the 'RF AM Stabiliser' board.

We also increased the power coming exiting the PSL table and going into the cavity from 11 microwatts to 20 microwatts by adjusting the polariser at the end of the table slightly.  The power has been set to 20 microwatts using the polariser a few days ago but had drifted down since then.

  7327   Fri Aug 31 10:24:36 2012 SteveUpdateVACchamber dog clamps checked

Quote:

I tightened as many of the dog clamps on the bottom of the BS, ITMX and ITMY chambers as I could find.  I used a torque wrench at 45 ft-lbs.  Some of the bolts of the dogs were too long, and I couldn't find an extender thing to accommodate the bolt so I could reach the nut.  None of the bolts moved that I was able to reach.

Steve, we're not doing final final alignment today (we will do it tomorrow), so please go around and double-check my work by checking all of the dogs first thing in the morning.  Thanks.

 Almost all chamber dog clamps on the floor checked. There are a few exception where it is impossible to to get to the nut. 

Only the OOC nuts turned little bit. So our elastomer discs are holding up well. This means that the chamber anchoring to the floor is good.

  7326   Fri Aug 31 10:16:02 2012 janosch, SteveUpdate ETMX, scattering preps

The alignment of the pick-off mirror near ETMX is done. Everything turned out to be easy once we realized that there is no sense getting the alignment laser (going through viewport to pick-off to ITMX) back to ETMX. It is only necessary to hit ITMX somehow, since this makes sure that there is one scattered beam that will make it from ITMX to pick-off through viewport.

After the auxiliary optic (that we never used in the end) was removed again, we levelled the optical table.

So in the current setup, we can have small-angle scattering measurements on ITMX and large-angle scattering measurements on ETMX.

Attachment 1: IMG_1609.JPG
IMG_1609.JPG
Attachment 2: IMG_1608b.jpg
IMG_1608b.jpg
  7325   Fri Aug 31 07:32:49 2012 SteveUpdateSUStarget for BS

Quote:

We installed beam targets on PRM and BS suspension cages.

On both suspensions one of the screw holes for the target actually houses the set screw for the side OSEM.  This means that the screw on one side of the target only goes in partial way.

The target installed on BS is wrong!  It has a center hole, instead of two 45 deg holes.  I forgot to remove it, but it will obvious it's wrong to the next person who tries to use it.  I believe we're supposed to have a correct target for BS, Steve?

The earthquake stop screws on PRM were too short and were preventing installation of the PRM target.  Therefore, in order to install the target on PRM I had to replace the earthquake stops with ones Jenne and I found in the bake lab clean room that were longer, but have little springs instead of viton inserts at the ends.  This is ok for now, but

WE NEED TO REMEMBER TO REPLACE EARTHQUAKE STOPS ON PRM WHEN WE CLOSE UP.

We checked the beam through PRM and it's a little high to the right (as viewed from behind).  Tomorrow we're going to open ITMX chamber so that we can get a closer look at the spot on PR2.

 The two eye  target for  the BS is in the clean tool box. It actually has irises.

  7324   Thu Aug 30 20:35:09 2012 jamieUpdateSUStarget installed on PRM, temporary earthquake stops in place

We installed beam targets on PRM and BS suspension cages.

On both suspensions one of the screw holes for the target actually houses the set screw for the side OSEM.  This means that the screw on one side of the target only goes in partial way.

The target installed on BS is wrong!  It has a center hole, instead of two 45 deg holes.  I forgot to remove it, but it will obvious it's wrong to the next person who tries to use it.  I believe we're supposed to have a correct target for BS, Steve?

The earthquake stop screws on PRM were too short and were preventing installation of the PRM target.  Therefore, in order to install the target on PRM I had to replace the earthquake stops with ones Jenne and I found in the bake lab clean room that were longer, but have little springs instead of viton inserts at the ends.  This is ok for now, but

WE NEED TO REMEMBER TO REPLACE EARTHQUAKE STOPS ON PRM WHEN WE CLOSE UP.

We checked the beam through PRM and it's a little high to the right (as viewed from behind).  Tomorrow we're going to open ITMX chamber so that we can get a closer look at the spot on PR2.

  7323   Thu Aug 30 20:31:35 2012 JenneUpdateIOOPZT1 and PZT2 set to center of their ranges

[Koji, Jenne]

Jamie and Koji pointed out that we need to be doing the in-vac alignment with the PZTs at the center of their ranges.  Also, we confirmed that they were set to "closed loop off", so the strain gauges were not supplying any feedback.

PZT1 was set to 0 for both pitch and yaw, since it has a very limited range of motion right now, so 0 is close enough.

For PZT2, Koji and I moved the slider in pitch and yaw, and watched the LCD output monitor on the PZT driver at the bottom of 1Y3.  We saw the value on the LCD change between slider values +4 to -6 for PZT2 yaw, so it is set to -1 as the center.  We saw the value on the LCD change between slider values -4 to +5 for PZT2 pitch, so it is set to +0.5 as the center.   Beyond these slider values (the sliders all go -10 to +10), the LCD value didn't change, either at 0, or at the maximum. 

Since PZT1 doesn't really move, this shouldn't affect any of the alignment work that Suresh and I did last night, although we should quickly confirm tomorrow. On the agenda for tomorrow is adjusting PZT2 such that we hit the center of PR2 (and hopefully that will also put us through the center of the PRM target, if the alignment was done well enough last time), so it's okay that we have only now set it to the center of its range.

  7322   Thu Aug 30 20:20:52 2012 JenneUpdateSUSWatek camera placed on SE viewport of ITMX to look at PRM

[EricQ, Jenne]

We placed the Watek camera on the SE viewport of the ITMX chamber, and focused it on the face of PRM.  We are not able to see any scattered light transmitted through the PRM, so this camera was an ineffective way to try to check spot centering on the PRM.  Jamie placed one of the new targets on the PRM cage - see his elog for details.

To get more use of the camera, we need to mount it on something, at the 5.5 inch beam height, and then cover that something with clean foil so we can place the camera on the table, in the beamline in various places.  We also need to carefully wrap the cables in foil so the don't dirty anything inside.

  7321   Thu Aug 30 19:08:08 2012 JenneUpdateVACDogs on BS, ITM chambers checked

I tightened as many of the dog clamps on the bottom of the BS, ITMX and ITMY chambers as I could find.  I used a torque wrench at 45 ft-lbs.  Some of the bolts of the dogs were too long, and I couldn't find an extender thing to accommodate the bolt so I could reach the nut.  None of the bolts moved that I was able to reach.

Steve, we're not doing final final alignment today (we will do it tomorrow), so please go around and double-check my work by checking all of the dogs first thing in the morning.  Thanks.

  7320   Thu Aug 30 18:01:05 2012 ElliUpdateIOOSetting up Input MC cavity scan measurement

Riju, Elli

Today tried to take our first cavity scan.  We unplugged the 55MHz sideband input from the RF combiner on the PSL table, and connected a network analyser instead.  Using the network analyzer we injected a 12dBm signal (swept from 32MHz to 45MHz) through the RF combiner into the EOM to create our swept sidebands.  We measured the  MC cavity response by looking at the signal comming out of the RF photodiode on the MC2 table.  I replaced the BNC cable connected to the RF PD with a longer BNC cable that could reach our network analyzer next to the PSL table.  Riju will post a diagram of our setup.

We didn't see the expected carrier resonances when we performed a cavity scan.  The light incident on the RF PD is around 0.7micro Watts and we are still thinking about whether this is strong enough to see our signal above the noise.  We also want to work out what the strength of our swept sidebands is.  We will attempt to do a 'real' cavity scan tomorrow.

 

  7319   Thu Aug 30 17:03:32 2012 janosch, Manasa, SteveUpdate ETMX

The baffle has been moved away from ETMX towards the edge of the table (in fact, it is a little beyond the edge). It is also rotated so that its long edge is horizontal. In this way it was possible to center the baffle hole with respect to the optical axis, but also make it possible that the camera looks over the baffle.

We have tried to get an alignment beam from view port -> ETMX pick-off ->ITMX-> back to EX. This work was pretty much unsuccessful though. We could see the green laser scattering around ITMX, but there was no good way to know when the beam hit ITMX. So tomorrow we will find a better way to check where the beam is hitting at ITMX and finish the alignment of the scattering pick-off mirror.

  7318   Thu Aug 30 13:10:41 2012 jamieUpdateCamerasETMX

Quote:

We have done some work at ETMX today. We installed the baffle and placed two mirrors on the table.

The baffle position/orientation still needs to be checked more thoroughly to make sure that the beam will pass through the center of the baffle hole.

I must say that I am not at all happy with the baffle situation.  It is currently completely blocking our camera view of the ETMX face.  Here's a video capture of the ETMX face camera:

etmx-face-baffle.png

The circle is the baffle hole, through which we can see just the bottom edge of the test mass.  I don't think whatever benefit the baffle gives out weights the benefit of being able to see the spot on the mirror.

This afternoon we will try to adjust the baffle, and maybe the camera view mirror, to see if we can get a better shot of the center of the TM.  If we can see the beam spot through the hole we can probably live with it.  If not, I think we should remove the baffle.

  7317   Thu Aug 30 12:01:27 2012 janosch, Manasa,SteveUpdate ETMX

We have done some work at ETMX today. We installed the baffle and placed two mirrors on the table.

The baffle position/orientation still needs to be checked more thoroughly to make sure that the beam will pass through the center of the baffle hole.

One of the two mirrors will stay on the table as pickoff. The other is only temporarily installed for alignment purposes. Later today we will shoot a laser into the chamber that will reflect off one of these mirrors towards the center of ITMX, then go back to the pickoff mirror next to ETMX and hopefully make it through the viewport.

To place the pickoff mirror, we had to move the "cable rack" next to ETMX a few inches towards the back of the table.

  7316   Thu Aug 30 08:37:11 2012 SteveUpdateSUSETMX and ITMX needs more attention

Quote:

ETMX appears to be fine.  It was stuck to its OSEMs in the usual way.  I touched it and it dislodged and is now swinging freely.  Damping loops have been re-engaged.

Screenshot.png

Earthquake m4.1 test for ETMX and moves ITMX.  ITMX-LR sensor 0.3V

Attachment 1: after4.1mAtm.png
after4.1mAtm.png
Attachment 2: eq4.1mAtm.png
eq4.1mAtm.png
  7315   Thu Aug 30 08:12:39 2012 SteveUpdateGeneralIn vac plans for tomorrow, 8/30

 

 1,PRM spot can be viewed directly from the window south-east of ITMX chamber.  I can easy set up the mobile- Watek for this reason or you can just use an IR viewer.

   Remember, we have 2 SOS centering targets ready to use , that Rana was suggesting.

2, PR2 spot centering can be viewed directly through window north-west of ITMX.

3, We should put back the BS view pick-up mirror for the vertical camera on the BS chamber and adjust its upper pick-up.

4, The BS centering can be viewed with the mobile-Watek placed inside the BS chamber immediately.

  7314   Thu Aug 30 00:08:34 2012 jamieUpdateGeneralIn vac plans for tomorrow, 8/30

Quote:

We need to check spot centering on PRM with camera tomorrow.

Suresh checked that we're not clipped by IP ANG/POS pickoff mirrors, but we haven't done any alignment of IP ANG/POS.

 I think we should NOT do any adjustment of IP ANG/POS now.  We should in fact try to recover them when doing the PRM spot centering

Quote:

Tomorrow:  Open ITMX door.  Check with Watek that we're hitting center of PRM.  Then look to see if we're hitting center of PR2.  Then, continue through the chain of optics.

The motivation for removing the ITMX door was so that the scatter measurement team could check alignment of the new viewing mirror next to ETMX.  After discussion today we decided that everything can be done at the X end.  They can inject a probe beam into the ETMX chamber, bounce it off of ITMX and align the viewing mirror with the reflection.  So we'll leave ITMX door on for now.

We should, however, inspect the situation ITMY and make sure we have good clearance in the Y arm part of the Michaelson.  Koji previously expressed suspicion that we might have clipping on the southern edge of the POY steering mirror, so we need to check that out.

Koji and I discussed the situation for getting camera face views of BS and PRM.  Koji said the original idea was to see if we could install something at the south-east view port of ITMX chamber.  Steve also suggested utilizing the "ceiling" camera mounted on the top of the IOO chamber.

Vertex tasks:

  • check spot centering in PRM
  • check that REFL is getting cleanly to the AP table
  • check IPPOS and IPANG - we should be adjusting IPPOS or IPANG at this point
  • check spot centering on BS
  • remove ITMY north door
  • check clearance of POY steering mirror
  • ...

in parallel:

  • Steve will inspect the situation for getting a camera view of BS and PRM face, either through IOO or ITMX.

End X tasks:

  • install baffle
  • install "permanent" ITMX viewing mirror, on west side of ETMX - this might require moving ETMX SUS cable bracket south
  • install temporary steering mirror for probe laser on south-east side of ETMX
  • at some point the scatter guys can also do transmission measurements of the ETMX view ports
  • ...
  7313   Wed Aug 29 21:02:45 2012 JenneUpdateIOOOptics between Faraday and PRM are centered, realigned

[Jenne, Suresh, with support from Jamie and Koji]

MC spots measured, MC1, MC3 no change.

No clipping going through Faraday.

Beam hitting to the right of center of PZT1.  It was translated sideways so we are now hitting it on the center.  Knobs adjusted so we hit center of MMT1.

Beam totally obscured by Faraday on the way to MMT2.  MMT2 moved north, so that we clear the Faraday by more than a beam diameter.  MMT1 knobs adjusted to hit center of MMT2.

MMT2 knobs adjusted to hit center of PZT2.

PZT2 didn't have enough range with knobs, so we loosened it, pointed then adjusted with knobs so we're hitting center of PRM. 

We need to check spot centering on PRM with camera tomorrow.

Suresh checked that we're not clipped by IP ANG/POS pickoff mirrors, but we haven't done any alignment of IP ANG/POS.

 

Tomorrow:  Open ITMX door.  Check with Watek that we're hitting center of PRM.  Then look to see if we're hitting center of PR2.  Then, continue through the chain of optics.

  7312   Wed Aug 29 20:43:23 2012 KojiUpdateIOOSetup for a cavity scan or the input mode cleaner

The technique is based on detection of the beating between the resonant carrier and a resonant higher order mode.
This means that the beat signal is cancelled out if the transmitted beam is integrated over the entire beam.
Thus, you may want to introduce intentional clipping (or cutting a half of the beam with a razor blade). 

Ref: LIGO DCC G080467: Precise Measurements on Longitudinal and Transverse Mode Spacings of an Optical Cavity using an Auxiliary Laser

I am quite curious on the measurement as I am going to do the same measurement for the aLIGO OMC.
I am interested in seeing the statistical evaluation on the precision of the measurement.

  7311   Wed Aug 29 19:28:41 2012 Elli KingUpdateLSCSetup for a cavity scan or the input mode cleaner

 Riju, Elli

Today we prepared our experimental setup to take a cavity scan of the input mode cleaner, which we want to measure in the next day or so.  Attached is a diagram of our setup.

What we want to do is to inject a set of sidebands into the PSL and sweep their frequency from 32-45 MHz (a range just over one fsr of the mode cleaner- vfsr=11MHz).  We will measure the power transmitted out of the MC using a photo-diode and demodulate this signal with our input signal from the Marconi.  From this we should be able to see the resonant frequencies of the carrier and the higher order modes.

One aspect we spent some time thinking about; whether we would be able to inject a signal into an EOM given the EOM and the Marconi are not perfectly impedance matched.  Based on Kiwamu’s previous e-log entries designing the EOM, we decided that injecting a signal in 32-45 MHz region at 15dBm is similar to injecting the 29.5MHz sideband (at the same power level with very similar input impedance.) Fingers crossed we don’t blow anything up first week on the job.

Attachment 1: 40m_cavity_scan_diagram.jpg
40m_cavity_scan_diagram.jpg
  7310   Wed Aug 29 17:35:34 2012 KojiUpdatePSLPower reduced!

The biggest reason why we could not lock the MC was that the beam was not properly hitting the MC REFL diode.

Now the MC REFL DC is about ~0.1 and 1.2 when the MC is and is not locked.

We increased the power according to the quantitative analysis of the intracavity power in this earlier entry

Autolocker script for the low power MC was modified so that the initial VCO gain is 3 in stead of 10.
The 2 steps of super boost were also enabled again.

  7309   Wed Aug 29 17:09:57 2012 jamieUpdateSUSETMX OK, free swinging

ETMX appears to be fine.  It was stuck to its OSEMs in the usual way.  I touched it and it dislodged and is now swinging freely.  Damping loops have been re-engaged.

Screenshot.png

  7308   Wed Aug 29 17:02:41 2012 ericqUpdatePSLPower reduced!

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 The power has been increased to 20mW. We got the 10mW number from the linked elog entry above. However, after venting we were having problems locking the MC. Upon investigating past elog posts, we found that 20mW was actually the power used in the past. The MC will now autolock. 

  7307   Wed Aug 29 12:46:49 2012 SteveUpdateVACVent completed in 4 hrs

 

ANTS ALERT please watch out for ants. We have them in the control room.

 

 

Attachment 1: vent72.png
vent72.png
Attachment 2: beforeVentOpl.png
beforeVentOpl.png
Attachment 3: beforeVent.png
beforeVent.png
  7306   Wed Aug 29 11:47:21 2012 ericqUpdateVACVenting

 [Steve, Eric]

I've been helping Steve vent this morning. The following things were done (from Steve's logbook):

  • Particle counts: 0.5 micron particles, 4200 counts per cubic ft
  • Vertex crane drive checked to be ok
  • Optical Levers set for local damping only
  • Saved some screens
  • PSL shutter and green shutters closed
  • HV Off checked, JAM nuts checked
  • Vac: Close V1, VM1, ans - VA6, open VM3 - RGA, cond: chamber open mode
  • 8AM: VV1 open to N2, regulator set  to 14 psi
  • 8:23AM: 35psi Instrument grade Air

(At this point, I took over the air canisters, while Steve made preparations around the lab. 

  • 9:00AM: 2nd air cylinder, 14 psi 
  • 9:40AM: 3rd air cyl
  • 10:20AM: 4th air cyl
  • 11:00AM: 5th air cyl

With the 5th cylinder, we began approaching 1 atm, so we slowed the regulator down to 5psi. Around 750 torr, Steve opened VV1 to air.

According to Steve, we will be at atmospheric pressure at  ~12:30pm.

  7305   Wed Aug 29 09:35:03 2012 janoschMetaphysicsRingdownripples 2

Ok, so the whole idea that mirror motion can explain the ripples is nonsense. At least, when you think off the ringdown with "pump off". The phase shifts that I tried to estimate from longitudinal and tilt mirror motion are defined against a non-existing reference. So I guess that I have to click on the link that Koji posted...

Just to mention, for the tilt phase shift (yes, there is one, but the exact expression has two more factors in the equation I posted), it does not matter, which mirror tilts. So even for a lower bound on the ripple time, my equation was incorrect. It should have the sum over all three initial tilt angles not only the two "shooting into the long arms" of the MC.

Quote:

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

 

  7304   Tue Aug 28 20:23:54 2012 KojiMetaphysicsRingdownripples

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

  7303   Tue Aug 28 19:21:37 2012 janoschMetaphysicsRingdownripples

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

Quote:

Isn't it just a ringing of the intracavity power as you shifted the laser frequency abruptly?

Quote:

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

 

 

  7302   Tue Aug 28 19:06:32 2012 KojiMetaphysicsRingdownripples

Isn't it just a ringing of the intracavity power as you shifted the laser frequency abruptly?

Quote:

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

 

  7301   Tue Aug 28 18:28:21 2012 janoschMetaphysicsRingdownripples

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

  7300   Tue Aug 28 17:59:03 2012 JenneUpdateVACVENT: GO!

I have turned of the high voltage supplies for PZT1 and PZT2.  The OMC PZT high voltage supplies were already off, since we aren't really using them currently.

I have closed the PSL shutter, but have not put in a manual extra beam dump yet.

All systems go for vent!

Steve - EricQ will be here around 8am to help with the vent.

  7299   Tue Aug 28 17:51:39 2012 JenneUpdatePSLPBS and Half Wave plates introduced

Quote:

[Jenne, Eric]

We installed a Half Wave Plate -> Polarized Beam Splitter -> Half Wave Plate in the PSL beam line, immediately after the EOM, to be used for attenuating the beam when we vent, as in Entry 6892.

It was illuminating to discover that the optics labeled QWP0-1064-10-2 are indeed half wave plates, instead of quarter wave plates as QWP suggests. 

The PBS transmits "P"/Horizontal polarization, but the beam coming from the EOM is "S"/Vertically polarized, and we want to keep that, since we do not want the beam attenuated quite yet. 

So, we use the HWP to rotate the P from the EOM to S, so that the majority of the power passes through the PBS. The second HWP then rotates the transmitted S back into P, which continues to the mode cleaner. When we want to attenuate, we will simply rotate the first HWP to change the proportion of S polarized light that will pass straight through the PBS and towards the mode cleaner. 

After setting the proper HWP angles, we aligned the PBS via minimizing the MC reflection.

Since we have not yet attenuated the power, we have not yet changed the BS for the MC reflection, since this would damage the PD. The beam splitter will be changed out for a 100% reflectivity mirror to increase the power to the PD when we do.

 

 Before we did this, I centered PSL POS and ANG, which gives us a reference of where the PSL beam was good when the MC spots were ~centered.  There had been a beam dump blocking them, possibly from the last time we put in the power attenuator optics.  This beam dump was moved a little to be out of the way of the PSL QPDs, and the PBS placed closer to the lens after the EOM, so that the PBS reflected beam is dumped.  However, we should not remove that razor dump when we remove the attenuation optics, since it is also dumping a stray IR beam from the PSL QPD pickoff windowd.

  7298   Tue Aug 28 17:43:04 2012 JenneUpdatePSLPower reduced!

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 I stopped the regular MC autolocker and told the crontab to startup the low power Mc autolocker on op340m.  Also, since we now have the new MC2 transmission setup, the power that gets to the 'regular' MC trans PD is lower, so I've lowered the lock threshold to 50 counts, from 100 counts.

  7297   Tue Aug 28 17:16:54 2012 ericqUpdatePSLPower reduced!

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

  7296   Tue Aug 28 17:02:16 2012 jamieUpdateGeneralsvn commit changes

I just spent the last hour checking in a bunch of uncommitted changes to stuff in the SVN.  We need to be MUCH BETTER about this.  We must commit changes after we make them.  When multiple changes get mixed together there's no way to recover from one bad one.

  7295   Tue Aug 28 16:27:22 2012 ericqUpdatePSLPBS and Half Wave plates introduced

[Jenne, Eric]

We installed a Half Wave Plate -> Polarized Beam Splitter -> Half Wave Plate in the PSL beam line, immediately after the EOM, to be used for attenuating the beam when we vent, as in Entry 6892.

It was illuminating to discover that the optics labeled QWP0-1064-10-2 are indeed half wave plates, instead of quarter wave plates as QWP suggests. 

The PBS transmits "P"/Horizontal polarization, but the beam coming from the EOM is "S"/Vertically polarized, and we want to keep that, since we do not want the beam attenuated quite yet. 

So, we use the HWP to rotate the P from the EOM to S, so that the majority of the power passes through the PBS. The second HWP then rotates the transmitted S back into P, which continues to the mode cleaner. When we want to attenuate, we will simply rotate the first HWP to change the proportion of S polarized light that will pass straight through the PBS and towards the mode cleaner. 

After setting the proper HWP angles, we aligned the PBS via minimizing the MC reflection.

Since we have not yet attenuated the power, we have not yet changed the BS for the MC reflection, since this would damage the PD. The beam splitter will be changed out for a 100% reflectivity mirror to increase the power to the PD when we do.

 

  7294   Tue Aug 28 11:28:31 2012 ericqUpdatePSLPMC alignment going bad

Quote:

PMC transmission started going down this afternoon, around 3pm-ish.  Right now it's 0.775, which is very, very low.  The new MC locking stuff is engaged, so it's not the FSS slow servo's fault. 

EDIT: I just realized that the limit of 0 counts output of the MC2 MCL filter bank was still engaged, from a time earlier this afternoon when I had switched back to the old servo, so there was no feedback going back to keep the slow drift of the laser in check.  PMC trans isn't coming back instantly, so I'll check it again when I come in tomorrow.

 

By adjusting the PMC steering mirrors, Jenne and I realigned the PMC input beam. Transmission is at 0.829 now. 

  7293   Tue Aug 28 09:37:33 2012 SteveUpdateVACRGA scan at day 56

RGA scan,  Maglev pumping speed at day 56

CC1 is dying. CC4 is real.

 

Attachment 1: pd72m56d.png
pd72m56d.png
Attachment 2: pd72m56dpr.png
pd72m56dpr.png
  7292   Tue Aug 28 00:23:54 2012 JenneUpdatePSLPMC alignment going bad

PMC transmission started going down this afternoon, around 3pm-ish.  Right now it's 0.775, which is very, very low.  The new MC locking stuff is engaged, so it's not the FSS slow servo's fault. 

EDIT: I just realized that the limit of 0 counts output of the MC2 MCL filter bank was still engaged, from a time earlier this afternoon when I had switched back to the old servo, so there was no feedback going back to keep the slow drift of the laser in check.  PMC trans isn't coming back instantly, so I'll check it again when I come in tomorrow.

Attachment 1: PMC_transmission_GoingDown_27Aug2012.png
PMC_transmission_GoingDown_27Aug2012.png
  7291   Tue Aug 28 00:16:19 2012 jamieUpdateGeneralAlignment and vent prep

I think we (Jenne, Jamie) are going to leave things for the night to give ourselves more time to prep for the vent tomorrow.

We still need to put in the PSL output beam attenuator, and then redo the MC alignment.

The AS spot is also indicating that we're clipping somewhere (see below).  We need to align things in the vertex and then check the centerings on the AP table.

So I think we're back on track and should be ready to vent by the end of the day tomorrow.

Attachment 1: as1.png
as1.png
  7290   Mon Aug 27 23:52:59 2012 JenneUpdateIOOMC Spots centered

Finally!

Jamie and I have the MC spots centered.  We did the normal move the input beam, realign jazz for a while, then when we got close, used the "move MC2 spot" scripts to get the MC2 spot back to ~center.

This was way easier when the measurements were real, and not just noise.  Funny that.

The dark blue spot is the farthest from 0 in pitch, and it is 1.04mm.  The cyan and yellow we've done a pretty good job of getting them equally far from zero.  Since we aren't translating the beam, we can't get better than the point at which the cyan and yellow curves cross.

Attachment 1: MCdecenter_26Aug2012.png
MCdecenter_26Aug2012.png
  7289   Mon Aug 27 18:59:24 2012 jamieUpdateIOOMC ASC screen was confusing - Jenne is not stupid

Quote:

We have figured out that some of these measurements, those with the WFS off, were also not allowing the dither lines through, so no dither, so no actual measurement.

Jamie is fixing up the model so we can force the WFS to stay off, but allow the dither lines to go through.  He'll elog things later.

In the c1ioo model there were filter modules at the output of the WFS output matrix, right before going to the MC SUS ASCs but right after the dither line inputs, that were not exposed in the C1IOO_WFS_OVERVIEW screen (bad!).  I switched the order of these modules and the dither sums, so these output filters are now before the dither inputs.  This will allow us to turn off all the WFS feedback while still allowing the dither lines.

I updated the medm screens as well (see attached images).

Attachment 1: Screenshot-1.png
Screenshot-1.png
Attachment 2: Screenshot-2.png
Screenshot-2.png
  7288   Mon Aug 27 18:32:48 2012 JenneUpdateIOOMC spot position - Jenne is stupid

Quote:

The MC REFL path was checked. ==> Some clippings were fixed. MC WFS is working now.

- MC was aligned manually

- The steering mirror for the WFS and camera was clipping the beam. => FIxed

- The WFS spots were realigned.

- There was small clipping on the MC REFL RFPD. ==> Fixed

 We have figured out that some of these measurements, those with the WFS off, were also not allowing the dither lines through, so no dither, so no actual measurement.

Jamie is fixing up the model so we can force the WFS to stay off, but allow the dither lines to go through.  He'll elog things later.

  7287   Mon Aug 27 17:14:00 2012 jamieUpdateCDSc1oaf problem

Quote:

I came in to the lab in the evening and found c1lsc had "red" for FB connection.
I restarted c1lsc models and it kept hung the machine everytime.

I decided to kill all of the model during the startup sequence right after the reboot.
Then run only c1x04 and c1lsc. It seems that c1oaf was the cause, but it wasn't clear.

The "red for FB connection" issue was probably a dead mx_stream on c1lsc.  That can usually be fixed by just restarting mx_stream.

There is definitely a problem with c1oaf, though.  It crashes immediately after attempting to start.  kernel log for a crash included below.

We will leave c1oaf off until we have time to debug.

[83752.505720] c1oaf: Send Computer Number  = 0
[83752.505720] c1oaf: entering the loop
[83752.505720] c1oaf: waiting to sync 19520
[83753.207372] c1oaf: Synched 701492
[83753.207372] general protection fault: 0000 [#2] SMP 
[83753.207372] last sysfs file: /sys/devices/pci0000:00/0000:00:1e.0/0000:2e:01.0/class
[83753.207372] CPU 4 
[83753.207372] Modules linked in: c1oaf c1ass c1sup c1lsp c1cal c1lsc c1x04 open_mx dis_irm dis_dx dis_kosif mbuf [last unloaded: c1oaf]
[83753.207372] 
[83753.207372] Pid: 0, comm: swapper Tainted: G      D    2.6.34.1 #5 X7DWU/X7DWU
[83753.207372] RIP: 0010:[<ffffffffa1bf7567>]  [<ffffffffa1bf7567>] T.2870+0x27/0xbf0 [c1oaf]
[83753.207372] RSP: 0000:ffff88023ecc1aa8  EFLAGS: 00010092
[83753.207372] RAX: ffff88023ecc1af8 RBX: ffff88023ecc1ae8 RCX: ffffffffa1c35e48
[83753.207372] RDX: 0000000000000000 RSI: 0000000000000020 RDI: ffffffffa1c21360
[83753.207372] RBP: ffff88023ecc1bb8 R08: 0000000000000000 R09: 0000000000175f60
[83753.207372] R10: 0000000000000000 R11: ffffffffa1c2a640 R12: ffff88023ecc1b38
[83753.207372] R13: ffffffffa1c2a640 R14: 0000000000007fff R15: 0000000000000000
[83753.207372] FS:  0000000000000000(0000) GS:ffff880001f00000(0000) knlGS:0000000000000000
[83753.207372] CS:  0010 DS: 0000 ES: 0000 CR0: 0000000080050033
[83753.207372] CR2: 000000000378a040 CR3: 0000000001a09000 CR4: 00000000000406e0
[83753.207372] DR0: 0000000000000000 DR1: 0000000000000000 DR2: 0000000000000000
[83753.207372] DR3: 0000000000000000 DR6: 00000000ffff0ff0 DR7: 0000000000000400
[83753.207372] Process swapper (pid: 0, threadinfo ffff88023ecc0000, task ffff88023ec7eae0)
[83753.207372] Stack:
[83753.207372]  ffff88023ecc1ab8 0000000000000096 0000000000000019 ffff88023ecc1b18
[83753.207372] <0> 0000000000014729 0000000000032a0c ffff880001e12d90 000000000000000a
[83753.207372] <0> ffff88023ecc1bb8 ffffffffa1c06cad ffff88023ecc1be8 000000000000000f
[83753.207372] Call Trace:
[83753.207372]  [<ffffffffa1c06cad>] ? filterModuleD+0xd6d/0xe40 [c1oaf]
[83753.207372]  [<ffffffffa1c07ae3>] feCode+0xd63/0x129b0 [c1oaf]
[83753.207372]  [<ffffffffa1c00dc6>] ? T.2888+0x1966/0x1f10 [c1oaf]
[83753.207372]  [<ffffffffa1c1b3bf>] fe_start+0x1c8f/0x3060 [c1oaf]
[83753.207372]  [<ffffffff8102ce57>] ? select_task_rq_fair+0x2c8/0x821
[83753.207372]  [<ffffffff8104cd8b>] ? enqueue_hrtimer+0x65/0x72
[83753.207372]  [<ffffffff8104d8f6>] ? __hrtimer_start_range_ns+0x2d6/0x2e8
[83753.207372]  [<ffffffff8104d91b>] ? hrtimer_start+0x13/0x15
[83753.207372]  [<ffffffff810173df>] play_dead_common+0x6e/0x70
[83753.207372]  [<ffffffff810173ea>] native_play_dead+0x9/0x20
[83753.207372]  [<ffffffff81001c38>] cpu_idle+0x46/0x8d
[83753.207372]  [<ffffffff814ec523>] start_secondary+0x192/0x196
[83753.207372] Code: 1f 44 00 00 55 66 0f 57 c0 48 89 e5 41 57 41 56 41 55 41 54 53 48 8d 9d 30 ff ff ff 48 8d 43 10 4c 8d 63 50 48 81 ec e8 00 00 00 <66> 0f 29 85 30 ff ff ff 48 89 85 18 ff ff ff 31 c0 48 8d 53 78 
[83753.207372] RIP  [<ffffffffa1bf7567>] T.2870+0x27/0xbf0 [c1oaf]
[83753.207372]  RSP <ffff88023ecc1aa8>
[83753.207372] ---[ end trace df3ef089d7e64971 ]---
[83753.207372] Kernel panic - not syncing: Attempted to kill the idle task!
[83753.207372] Pid: 0, comm: swapper Tainted: G      D    2.6.34.1 #5
[83753.207372] Call Trace:
[83753.207372]  [<ffffffff814ef6f4>] panic+0x73/0xe8
[83753.207372]  [<ffffffff81063c19>] ? crash_kexec+0xef/0xf9
[83753.207372]  [<ffffffff8103a386>] do_exit+0x6d/0x712
[83753.207372]  [<ffffffff81037311>] ? spin_unlock_irqrestore+0x9/0xb
[83753.207372]  [<ffffffff81037f1b>] ? kmsg_dump+0x115/0x12f
[83753.207372]  [<ffffffff81006583>] oops_end+0xb1/0xb9
[83753.207372]  [<ffffffff8100674e>] die+0x55/0x5e
[83753.207372]  [<ffffffff81004496>] do_general_protection+0x12a/0x132
[83753.207372]  [<ffffffff814f17af>] general_protection+0x1f/0x30
[83753.207372]  [<ffffffffa1bf7567>] ? T.2870+0x27/0xbf0 [c1oaf]
[83753.207372]  [<ffffffffa1c06cad>] ? filterModuleD+0xd6d/0xe40 [c1oaf]
[83753.207372]  [<ffffffffa1c07ae3>] feCode+0xd63/0x129b0 [c1oaf]
[83753.207372]  [<ffffffffa1c00dc6>] ? T.2888+0x1966/0x1f10 [c1oaf]
[83753.207372]  [<ffffffffa1c1b3bf>] fe_start+0x1c8f/0x3060 [c1oaf]
[83753.207372]  [<ffffffff8102ce57>] ? select_task_rq_fair+0x2c8/0x821
[83753.207372]  [<ffffffff8104cd8b>] ? enqueue_hrtimer+0x65/0x72
[83753.207372]  [<ffffffff8104d8f6>] ? __hrtimer_start_range_ns+0x2d6/0x2e8
[83753.207372]  [<ffffffff8104d91b>] ? hrtimer_start+0x13/0x15
[83753.207372]  [<ffffffff810173df>] play_dead_common+0x6e/0x70
[83753.207372]  [<ffffffff810173ea>] native_play_dead+0x9/0x20
[83753.207372]  [<ffffffff81001c38>] cpu_idle+0x46/0x8d
[83753.207372]  [<ffffffff814ec523>] start_secondary+0x192/0x196

  7286   Mon Aug 27 15:49:46 2012 JenneUpdateSUSEQ damage

Quote:

After shaking ITMX by the alignment bias in yaw, it came back.

As ETMX seems to be largely misaligned yaw (and did not come back with the alignment impact),
the condition of the magnets are not clear. Only the side OSEM is responding nicely.

Quote:

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

I tried to take some photos through the window of ETMX's chamber, to see if I could see any magnets.  What we have learned is that Jenne is still not the world's best photographer.  I was holding the camera at ~max zoom inside the beam tube between the table and the window, so that's my excuse for the photos being fuzzy.  The only thing that I can really conclude is that the magnets look like they are still there, but Jamie thinks they may be stuck on the PDs/LEDs (now looking at the photos myself, I agree, especially with UL and LR). 

It looks like the best thing to do at this point, since Koji already tried jerking ETMX around in yaw a little bit, is just wait and open the door, to see what's going on in there.  I posted the photos on Picasa:

https://picasaweb.google.com/foteee/ETMX_MaybeStuck_ThroughWindow_27Aug2012

I propose that, if the magnets are broken, we pull the ETM out of the camber and fix it up in the cleanroom while we pump back down.  This would restrict us from doing any Xarm work, but will force me to focus on DRMI, and we can put the ETM back when we vent to install the tip tilts.

  7285   Mon Aug 27 15:46:55 2012 SteveUpdateSAFETYsafety training

 Rijuparna Chakraborty and Elli Elenora King received 40m specific  basic safety training in the 40mLab

Attachment 1: IMG_1597.JPG
IMG_1597.JPG
  7284   Mon Aug 27 12:03:54 2012 KojiUpdateIOOMC spot position - callin' it quits

The MC REFL path was checked. ==> Some clippings were fixed. MC WFS is working now.

- MC was aligned manually

- The steering mirror for the WFS and camera was clipping the beam. => FIxed

- The WFS spots were realigned.

- There was small clipping on the MC REFL RFPD. ==> Fixed

  7283   Mon Aug 27 10:49:03 2012 KojiUpdateSUSEQ damage

After shaking ITMX by the alignment bias in yaw, it came back.

As ETMX seems to be largely misaligned yaw (and did not come back with the alignment impact),
the condition of the magnets are not clear. Only the side OSEM is responding nicely.

Quote:

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

  7282   Mon Aug 27 09:24:17 2012 SteveUpdateSUSEQ damage

  It looks like we may lost 1 (or 3 )  magnets? Do not panic, it's not for sure

 

Attachment 1: eqDamage.png
eqDamage.png
  7281   Mon Aug 27 08:34:18 2012 SteveUpdatePEMearthquakes

Shasky day yesterday postpones venting. We had about 11 shakes larger than mag 4.0 Mag5.5 was the largest at  13:58 Sunday, Aug 26 at  the Salton Sea area.

Atm3,  ITMX and ETMX  did not come back to it's position

Attachment 1: eq5.5Msaltonsea.png
eq5.5Msaltonsea.png
Attachment 2: M5.5inaction.png
M5.5inaction.png
Attachment 3: EQeffect.png
EQeffect.png
  7280   Mon Aug 27 01:05:36 2012 JenneUpdateIOOMC spot position - callin' it quits
spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[-0.98675603448324423, -0.94064212026141558, 2.6749179375892544, -0.65896393156684185, -0.4508281650731974, -0.55109088188064204]

MC3 pitch isn't what I'd like it to be, but MC1 and MC3 pitch aren't quite acting in relation to each other how I'd expect. Sometimes they move in common, sometimes differentially, which is confusing since I have only ever been touching (on the PSL table) the last steering mirror before the beam is launched into the vacuum.

The latest few measurements have all been with the WFS off, but reflection of ~0.48 . I haven't figured out why yet, but MC1 and MC3 yaw WFS outputs start to escalate shortly after the WFS becoming engaged, and they keep knocking the MC out of lock, so I'm leaving them off for now, to be investigated in the morning.
Attachment 1: MCdecenter_26Aug2012.png
MCdecenter_26Aug2012.png
  7279   Sun Aug 26 21:47:50 2012 KojiUpdateCDSC1LSC ooze

I came in to the lab in the evening and found c1lsc had "red" for FB connection.
I restarted c1lsc models and it kept hung the machine everytime.

I decided to kill all of the model during the startup sequence right after the reboot.
Then run only c1x04 and c1lsc. It seems that c1oaf was the cause, but it wasn't clear.

ELOG V3.1.3-